Review Correct A small wooden box is connected to a spring, and is allowed to oscillate horizontally on a flat surface. The box has mass 170 g, and the spring has spring constant 2.7 N/m The box moves with a velocity of 21 cm/s when zo =-5.2 cm ▼ Part C At the point where the boxs acceleration is maximum, what is the boxs position? Express your answer to two significant figures and include the appropriate units. cm Submit Previous Answers Request Answer X Incorrect; Try Again Part D What is the speed of the box when 3.0 cm? Express your answer to two significant figures and include the appropriate units. Value Units Submit Request Answer
Answer #1

v=w(A^2-x^2)^.5

v=velocity

w=angular frequency=(k/m)^.5=(2.7/.17)^.5=4rad/s

A=amplitude

from given data

21=4(A^2-5.2^2)^.5

27.76=A^2-5.2^2

A^2=54.8

A=7.4cm

Maximum acceleration at amplitude position i.e 7.4cm

Speed when x=3m

4(7.4^2-3^2)^.5

=27.07cm/s